OpenStudy (anonymous):

Ice Cream Man needs a lot of help... What is the domain and range of Sin^-1x? Ice Cream Man knows that the Domain is [-1,1] Ice Cream Man knows that the Range is -pi/2, pi/2. What Ice Cream Man does not know is that there are different kinds of brackets surrounding the range. One of the options is (-pi/2,pi/2) and the other option is [-pi/2,pi/2]. Ice Cream Man appreciates your help in advance.

OpenStudy (anonymous):

I like talking in 3rd person too XD

OpenStudy (anonymous):

Ice Cream Man does not know what you are talking about and only wants help Mr. John

OpenStudy (anonymous):

@mmmmmmmmmm123

OpenStudy (jdoe0001):

\(\large { \textit{Inverse Trigonometric Identities} \\ \quad \\ \begin{array}{cccl} Function&{\color{brown}{ Domain}}&{\color{blue}{ Range}}\\ \hline\\ {\color{blue}{ y}}=sin^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1&-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \\ \quad \\ {\color{blue}{ y}}=cos^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1& 0 \le {\color{blue}{ y}}\le \pi \\ \quad \\ {\color{blue}{ y}}=tan^{-1}({\color{brown}{ \theta}})&-\infty\le {\color{brown}{ \theta}} \le +\infty &-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \end{array} }\)

OpenStudy (anonymous):

ice cream man knows this already but what ice cream man needs to know is what kind of parenthsis to put, [ ] or ( )

OpenStudy (jdoe0001):

hehe well... look at the \(\Large \le\qquad \ge\) signs :) that'd say which is which

OpenStudy (anonymous):

ice cream man does not understand.... :(

OpenStudy (jdoe0001):

well... what does (this, that) mean? or what does [this, that) mean? or what does [this, that] mean?

OpenStudy (anonymous):

ice cream man will guess that the answer is ( )

OpenStudy (anonymous):

is ice cream man right? @jdoe0001

OpenStudy (jdoe0001):

hmm you may want to shake your 8ball then

OpenStudy (jdoe0001):

methiks icecreamman is slacking off in his interval notation

OpenStudy (anonymous):

ice cream man would now like to change the answer to [ ]

OpenStudy (anonymous):

John does not know the answer. John is sad

OpenStudy (anonymous):

Ice Cream Man is sorry that John is sad and would like to give John an ice cream cone

OpenStudy (jdoe0001):

methinks icecreamman needs to reread his section on interval notation

OpenStudy (anonymous):

John is eating ice cream right now. Greta sad because you won't share.

OpenStudy (anonymous):

Greta did not help. So Greta does not get an ice cream.

OpenStudy (anonymous):

It does not give the option of [ ) @jdoe0001

OpenStudy (anonymous):

But Greta wants ice cream. (for real though, John won't share his ice cream) :(

OpenStudy (anonymous):

Greta will get ice cream if she helps answer the question :)

OpenStudy (anonymous):

@Directrix @freckles will you please help ice cream man?

OpenStudy (anonymous):

@Nnesha ice cream man needs help will you help?

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

me thinks icecreamman needs to reread his section on interval notation

OpenStudy (anonymous):

OH SO ITS [ ] RIGHT?!?!?!?! @kidrah69

OpenStudy (kidrah69):

@jdoe0001